OpenStudy (anonymous):

Find the unit vector having the same direction as v. v = -3i + j

OpenStudy (anonymous):

unit vector having same direction as x is given by = \[\frac{ Vector \space x }{magnitude \space of \space Vector \space x}\]

OpenStudy (anonymous):

magnitude of your vector is \[\sqrt{3^2+1^2}=\sqrt{10}\]